trying to help my daughter with these two (Part II)
Matthew's grandmother left him $10,000 for college. Matthew puts some of the money into a savings account earning 3% interest per year. With the rest, he buys a certificate of deposit earning 5% per year. If matthew puts $300.00 in savings, how much money will he have in one year? (Write the Equation)
Thanks again,

Answer by mananth(16949)

$300 in savings account 3%
10,000-300 = $9,700 certificate 5%
Interest = p*r*t
p= principal= 300
r=>rate=>5%=0.05
t= 1 year
savings interest = 300*1*0.03=$9

Certificate interest = 9700*1*0.05=485
Total interest = 485+9 = $494
$10,000+$494= $10,494
Total amount at the end of the year = $10,494. 00